OpenStudy (moneybaby980):

Solve for x: - 4/ 3* x - 6 equals - 26 −27 −15 15 27

OpenStudy (anonymous):

okay so what would you do first ? \[-4/3*x-6=-26\]

OpenStudy (moneybaby980):

ok

OpenStudy (anonymous):

what would you do first ?

OpenStudy (anonymous):

personally i would subtract -6 from both sides you would end up with \[-4/3*x=-20\]

OpenStudy (moneybaby980):

so what happens to the x

OpenStudy (anonymous):

well you still have -4/3

OpenStudy (moneybaby980):

do I divide 4 by 3

OpenStudy (anonymous):

im pretty sure it would be that way but -4 instead of just 4

OpenStudy (moneybaby980):

oh ok I got 1.3

OpenStudy (anonymous):

did you use your negative sign cause i got -1.3

OpenStudy (anonymous):

but same difference what would you do next?

OpenStudy (moneybaby980):

I keep forgetting to put -

OpenStudy (anonymous):

so now the equation is \[-1.3*x=-20\]

OpenStudy (anonymous):

i just did the whole equation and the answer i got isnt one of your answer choice.... so your best bet is to get a second opinion or id say -27